I'd left Ash to climb the stairs to the ward and carried on to school with Bo in time take myself to the dentist by 9am. Ash and Ally went to a wedding in Somerset and I did a stint with Pads from about 10:30 At noon I started to journal it:
12:10 anti sickness and antibiotics go in/up
12.15 feed off - feeling ill
12.20 was ill - couldnt decide what position to get in as the preferred leaning over is too painful.
13.50 Consultant (and Moli) are there when I return to the bedside. Moli always gets the best out of him. They watch a Bond film together. Change of antibiotic to vancomycin. We're seeing nothing of the anticipated climb in blood counts, including via the special test about the stem cells. Just got to focus on getting him better. With the feed off, I've got to keep the water going in via a syrynge, 30mls slowly at a time.
14.30 Movicol going in 30mls at a time - oramorph given.
14.40 Vancomycin infused - plus 500 mls bolus infused (i.e. it's pumped through fast into his hickman line, to boost his blood pressure and force his circulation to push blood into his extremities) - cold 'corned beef' arms and legs has got the staff worrying he's going septic.
15.00 Nursing auzilliary is back to do obs: Temp is 36.9
15.20 Registrar came to check on P who had just had one. Consultant arrives too - they're a bit worried about him vasodilating
16.30 tazocin going in - nurse still thinks he's mottled and cool. Two SHO's arrive: asking and squeezing, probing and poking. Conclusion is to give more fluid IV 500mls 'bolus'.
16.40 I put the feed back on at 20mls/hr although pads was asking for 10. Ten?! what's the point of 10?
17.05 NA's back again: BP is 85/64 Temp 36.9
17.30 Reg is back - seemed happy enough
17.40 Obs done again: 91/63 37.2 I accelerated feed back up to 40mls/hour
17.55 Consultant's back - though everyone was gettin jumpy, Pads seemed to have improved a lot - not so blotchy and he is chirpy. Oramorph just given - at 5mls not 7.5mls but he can have more if needed for 'breakthrough pain'. Inflamation around mic-key button seems to have moved below it, not above. Pads is still not sure why everyone was interested and concerned about his 'mottled' skin. The consultant spells it out although I still have to explain the word 'septic' to him. They still think the mic-key is source even tho it didn't grow anything in the swab/cultures.
In other news, they can do the harvest next week if needed as they can't give chemo if he's so splatted anyway.
18.30 Platelets arrive as the count for them was 20 (ten below the threshold)
Gonna try pushin this last 60 mls of movicol through in a moment.
19.00 He watches a film clutching bowl to catch anything from his mouth, on the commode behind the curtain, the nurse is flushing his platelets, and the NA's taking his obs (37.3 83/59), with me standing over him and Auntie Sarah trying to take over so I can get Bo from his friend's house. Took Bo to the park and chilled a bit with him.
Ally and Ash return having had a good day. She goes in to relieve Sarah.
Spoke to Ally at 11.30pm - He's been having a lot of pain getting out of bed so they've given paracetamol - he can get more oramorph if needed too. Thankfully the little boy sharing the room has calmed down and now they're both asleep.
Latest obs show his Temp was 37.6 with the feed still going through at 10mls.
He was really getting buzzed by the staff today. They were quite worried about his condition. Next stop is potentially ICU, but he was protesting that his hands are often mottled and cold even when he's well!